### 08 Dec 2023 v3

# Evaluating the Ebb and Flow: An In-depth Analysis of Question-Answering Trends across Diverse Platforms

Abstract
Community Question Answering (CQA) platforms steadily gain popularity as they provide users with fast responses to their queries. The swiftness of these responses is contingent on a mixture of query-specific and user-related elements. This paper scrutinizes these contributing factors within the context of six highly popular CQA platforms, identified through their standout “answering speed”. Our investigation reveals a correlation between the time taken to yield the first response to a question and several variables: the metadata, the formulation of the questions, and the level of interaction among users. Additionally, by employing conventional machine learning models to analyze these metadata and patterns of user interaction, we endeavor to predict which queries will receive their initial responses promptly.
Keywords:
community question answering response time asker-answerer graph classification model
1 Introduction
Community question-answering platforms are gradually evolving over the past decade. Such portals allow users to post their queries and answer the questions. These CQA platforms help experts share their knowledge and new users solve their queries. Initially, these community question-answering platforms focused on providing the most relevant answers to the users’ queries [14][2]. As the CQA platforms keep growing, they not only focus on providing the relevant answers but also on quickly delivering such answers [12][11]. The growth and maturity of such platforms is contingent on a number of variables, including the “quality of questions and answers”, the “response time of queries”, “user involvement”, and “user activity” [8]. The users with expertise in specific domains significantly help the platform’s growth. Based on the users’ activeness on the platform and the importance of their posts, they have been rewarded reputation scores, badges, and additional privileges. In most such CQA platforms driven by a reputation system, the platform’s popularity depends on the response time [3] of the question (i.e., time to receive the first answer). So, the users who want to post a query may also want to know the time by which they can expect a response to their question. StackExchange [^1] https://stackexchange.com/ is one of the CQA platforms where the users can ask questions and receive answers from other users. An analysis of such a platform’s dynamics helps the community maintainers adapt and design the framework to address various CQA-related problems [1] [7] [13] [9] [6] [5] [4]. In this paper, we examine the dynamics of different CQA platforms relating the usual response time with other factors of the portals. We conduct an empirical study on six CQA platforms – Mathematics, Software Engineering, English, Game Development, Chemistry, and Money. To analyze the dynamics of the diverse community question-answering platforms, we look into various factors mentioned here – (1) Characterize the questions in terms of textual information, linguistic characteristics, question tags, votes received by the questions etc. We obtain correlations of these features with the time required to respond to a question. (2) Characterizing the users of the platform in terms of their reputation & activity in the platforms. Further, we build Asker-Answerer graph (AAG) where the nodes are the users and a link is created if the question from a user is answered by another user. Subsequently, we conduct in depth analysis of the characteristics of the common users present across all the platforms. (3)For all the platforms, we use metadata of the questions, question structure and user interaction as features to build a standard machine learning model for classifying the newly posted questions which are having fast answer. The majority of preceding studies [10] scrutinize the correlation between diverse question-related factors and response time. The novelty of our current research resides in our endeavor to comprehend how various inherent network-related features (reflected in the asker-answerer graph) coincide with question-related features in relation to response time. In this study, we make the following observations – (1) Short questions get faster responses. Chemistry and Software Engineering questions are harder. More tags mean slower replies. Math, Software Engineering, and English answers come in an hour, others in a day. In Software Engineering and English, many questions get $\geq 3$ answers.(2) Most users ( 60%-70%) on all platforms stay inactive. Domains like Money, Game Development, and Chemistry have more active users than Mathematics, Software Engineering, and English. (3) XGBoost and MLP perform best at classifying new questions on most platforms. The second best is usually the random forest model.
2 Dataset
We test six CQA datasets from StackExchange up to March 2022, picking six platforms (top three and bottom three in answering speed) from an initial group that includes Physics, Chemistry, Mathematics, English, Software Engineering, Game Development, AskUbuntu, Mathematica, Travel, Money. Answering speed, different from response time, is the % of questions answered within time $t$ ($t$ = 10 mins, 20 mins, etc.). Math, English, and Software Engineering are top in speed; Money and Finance, GameDev, and Chemistry are bottom. Data includes question title, body, tags, reporting time, answers with timestamps, votes, and user reputations. Table 1 presents dataset stats.
Datasets #Questions #Unique Tags #Unique Users Age (in yr.)
SE 60,801 1667 3,46,642 12 (2010)
GD 53,427 1094 1,22,553 12 (2010)
MA 14,79,363 1898 8,88,141 12 (2010)
EN 1,24,373 981 3,50,766 12 (2010)
CH 40,742 368 88,677 10 (2012)
MO 35,494 1005 84,650 13 (2009)
Table 1: Basic statistics of six community question answering platforms.
Datasets Average number Total number % of answered % of questions Average number
of tags of answers questions having accepted answer of votes
SE 2.802 1,72,275 95.44 57.78 6.6139
GD 2.795 77,996 86.31 52.80 2.567
MA 2.369 19,76,51 82.32 52.71 2.11
EN 2.08 2,79,346 92.12 48.1 3.36
CH 2.396 47,420 79.6 40.63 3.42
MO 3.11 66,917 91.19 45.32 4.71
Table 2: Basic statistics about the question structure
3 Empirical analysis
In order to compare the platforms we consider two different factors – (i) the characteristic features of the questions posted in a platform, and (ii) the interaction behaviour of users on a platform. In each case we correlate these factors with the response time, which is the time required to obtain the first answer after a question is posted.
3.1 Characterizing the questions
In order to characterize the questions of different platforms, we explore various structural properties of the question. These include length of the title, length of the body, linguistic characteristics of question title and question body. In Table 2, we have shown the average number of tags present in the question, total number of answers, percentage of questions having at least one answer, percentage of questions having accepted answer, average number of votes received by questions. From Table 2, it is observed that 90%-95% of questions have received at least one answer in $SE$, $EN$ and $MO$ platforms. Lowest percentage ($\sim$79%) of questions have been answered in the $CH$ platform. Further $SE$, $GD$ and $MA$ have a higher percentage of questions having accepted answers than other platforms. The average number of votes to the questions are maximum for the $SE$ platform.
Length of the title and body: We have computed length of the title and body as the number of words present in the each of them respectively. Further we took top 20% questions ($Q_{top}$) and bottom 20% questions ($Q_{bottom}$) based on the response time. Top 20% questions correspond to a lower response time and bottom 20% questions correspond to a higher response time. In $MA$, $EN$ and $SE$, the average length of the body of $Q_{top}$ questions are 69.06, 71.86, 150.70 respectively; the length of the body of $Q_{bottom}$ questions are 117.48, 102.55, 189.70 respectively. Thus questions with lesser length get faster answers. The same trend is observed for $MO$, $GD$ and $CH$ where the length of the body for $Q_{top}$ are 121.56, 133.63, 81.197 and for $Q_{bottom}$ are 138.48, 166.15, 106.117 respectively. We did not find much difference in the length of the titles for the $Q_{top}$ vs the $Q_{bottom}$ questions.
Linguistic characteristics: In order to understand the linguistic characteristics, we have used three measures – Flesch reading ease test measure  [^2] https://simple.wikipedia.org/wiki/Flesch_Reading_Ease (higher the better), Coleman Liau Index  [^3] https://en.wikipedia.org/wiki/Coleman-Liau_index (lower the better) and Automated Readability Index  [^4] https://en.wikipedia.org/wiki/Automated_readability_index(lower the better). We observe the readability scores for $Q_{top}$ and $Q_{bottom}$ questions. For Flesch-reading-ease test (see Figure 1), the $Q_{top}$ questions of $MA$, $EN$ and $SE$ have average scores of 68.53, 71.71, 63.31 respectively while the $Q_{bottom}$ questions have the average scores of 60.59, 68.86 and 58.14 respectively. In case of $MO$, $GD$ and $CH$, the average scores of $Q_{top}$ are 71.42, 62.75 and 63.22 respectively while that of $Q_{bottom}$ questions are 68.83, 55.26 and 58.87 respectively. Thus readability is better (easier) for the $Q_{top}$ questions across all platforms irrespective of their popularity (in terms of the answering speed). This observation holds true for the other two readability metrics as well (see Figure 1).
Figure 1: (left): a. Flesch Reading Ease, ($2^{nd}$ from left): b. Coleman-Liau index, (middle): c. Automated readability index scores for $Q_{top}$ and $Q_{bottom}$ questions. ($2^{nd}$ from right): d. Average Clustering Coefficient, (right): e. Average Degree of AAG of $Q_{top}$ and $Q_{bottom}$.
Tag related analysis: For each platform, we obtain the most frequently occurring tags (top 50) across the questions posted. For $SE$, Java, Design, C# are the most popular tags. For $GD$, Unity, C++, C# and OpenGL are most frequent and for $MO$, United States and taxes are most popular tags. In case of $MA$, $EN$ and $CH$, we do not observe any particularly dominant set of tags. Further, for each portal, we calculate the average number of tags present in $Q_{top}$ and $Q_{bottom}$ questions. In $MA$, $EN$ and $SE$, the average number of tags in $Q_{top}$ and $Q_{bottom}$ are 2.08, 2.03, 2.53, 2.66, 2.17 and 3.02 respectively. In $GD$, $CH$ and $MO$, the average number of tags for $Q_{top}$ and $Q_{bottom}$ are 2.58, 2.28, 3.00, 2.94, 2.54 and 3.23 respectively. We observe that there are lesser number of tags in the questions in $Q_{top}$ than in $Q_{bottom}$. This could possibly indicate lesser tags makes a question topically less confusing and streamlines it more effectively to people who can actually answer them.
Votes received by the question: In order to understand the temporal pattern of votes received by the questions, we count the fraction of questions receiving 1 upvote, 2 upvotes and $\geq 3$ upvotes on three consecutive days, i.e., the same day when the question is posted followed by the two subsequent days. In the same day, $\sim$30questions of $SE$ platform receive more than 3 votes. But for other platforms most of the questions received one vote. In second day, around $\sim$20of questions receive one more vote in $SE$, $MO$ and $CH$ platforms. In case of $SE$ and $MO$, there are 10% of questions receiving at least three votes in the second day. In the third day, there are very less number of questions receiving votes across all the platform. This indicates that on all platforms the attention span of users for a particular question dies down within 48 hours from the post of the question. In case of $MA$, $EN$ and $SE$, the average votes for the questions in $Q_{top}$ are 1.99, 5.31 and 15.82 respectively and that for the questions in $Q_{bottom}$ are 2.75, 2.60 and 4.44 respectively. In case of $MO$, $GD$ and $CH$, the average number of votes in $Q_{top}$ are 6.46, 4.63 and 4.21 respectively while that in $Q_{bottom}$ are 2.96, 2.08 and 5.17 respectively. Hence questions in lower response time typically receive more upvotes irrespective of the popularity of the platform.
3.2 Characterizing the users
In this section we analyse the user behaviour on each of these platforms. Number of users in $MA$, $SE$ and $EN$ are approximately  [^8] K, 346K and 350K which is much larger than the number of users in other platforms. For $MO$, $GD$ and $CH$, the number of users are 84K, 122K and 88K respectively. In Table 3, we show the basic statistics about these unique users such as % of users only posting question, % of users only answering questions, % of users posting both the question and answers, and users whose at least one answer got accepted by an asker.
Datasets %Users posting only questions %Users posting only answers %Users posting both Q&A %Users whose at least one answer got accepted
SE 8.9 7.5 2 1.9
GD 19 12.7 5.3 4.9
MA 33 10 5 4
EN 17 12.3 2.6 2.3
CH 20 7 2.6 2.1
MO 22.2 9.8 2.7 2.2
Table 3: Basic statistics of the user interaction.
We have seen that there are around 60%-70% users who remain inactive in all the platforms. Inactive users neither post anything, nor do they interact with anyone on the platform. $MA$, $SE$ and $EN$ platforms have approximately 37.38%, 14.14% and 26.73% of users who have posted at least one question or one answer (aka active users). In case of $MO$, $GD$ and $CH$, the percentage of active users are 29.38%, 26.47% and 25.29% respectively. According to Table 3, $MA$ has maximum fraction of users (33% users) who have posted at least one question. In terms of writing answers, around 12% users are active in $EN$ and $GD$ platforms which is larger than the other platforms. $GD$ has maximum fraction of users ($\sim$5.3%) who have posted both question and answer. Once again, $GD$ has the largest fraction of users ( 4.9%) whose at least one answer got selected. Further we conduct in depth analysis on the users who have asked or answered any questions in $Q_{top}$ and $Q_{bottom}$ respectively. In Table 5, we show number of askers, answerers and their overlap for $Q_{top}$ and $Q_{bottom}$ users. We analyze various factors of the asker and answerer of the questions in $Q_{top}$ and $Q_{bottom}$ below.
Reputation score of users: We separately observe the reputation scores of asker and answerer of questions of $Q_{top}$ and $Q_{bottom}$ category. Asker is the user who asked the question in the portal and answerer is the user who answered the question. In $MA$, $EN$ and $SE$, the average reputation scores of askers in $Q_{top}$ are 559.07, 546.27 and 1001.96 respectively. The average reputation scores of askers in $Q_{bottom}$ 705.52, 522.78 and 644.14 respectively. In $MO$, $GD$ and $CH$, the average reputation score of askers in $Q_{top}$ are 585.71, 401.53 and 424.68 respectively. The average reputation score of askers in $Q_{bottom}$ are 540.71, 307.14 and 561.10 respectively. In case of answerers, the average reputations in $Q_{top}$ are 4063.39, 2147.42 and 3209.59 for $MA$, $EN$ and $SE$ respectively and those in $Q_{bottom}$ are 2001.24, 1622.79 and 1846.81 respectively. For $MO$, $GD$ and $CH$, the average reputation of answerer in $Q_{top}$ are 4017.05, 1117.52 and 1948.33 respectively while that in $Q_{bottom}$ are 1965.42, 552.84 and 1397.38 respectively. We observe three things here – (i) across all platforms the reputation scores of askers as well as answerers are typically more for $Q_{top}$ questions (ii) in the popular platforms the reputation scores for both askers and answerers are higher and (iii) answerer reputations are an order of magnitude higher than the askers.
Asker-Answerer graph (AAG): To understand the user interaction patterns on a platform, we build an undirected graph where the nodes are the users and the edges denote if a user (answerer) answers a question posted by another user (asker). For every platform, we build two such AAGs for users who either asked or answered the questions of $Q_{top}$ and $Q_{bottom}$ respectively (See Table 4). Further, we calculate average degree, average clustering coefficient for the whole network. In Figure 1 (e), we observe that the average degree of $Q_{top}$ AAG for all platforms except $CH$ are relatively higher than the average degree of $Q_{bottom}$ AAG. For all the platforms, the average clustering coefficient of $Q_{top}$ AAG are higher than the average clustering coefficient of $Q_{bottom}$ AAG (see Figure 1 (d)). This indicates that there is a higher density of interactions among the asker-answer(s) in the $Q_{top}$ AAG (and exceptionally higher platforms $SE$, $GD$ and $MO$).
Datasets # nodes # edges fraction of nodes present in largest connected component.
SE ($Q_{top},Q_{bottom}$) 10319, 12756 11179, 11234 0.73, 0.53
GD ($Q_{top},Q_{bottom}$) 8011, 9453 8730, 8547 0.72, 0.42
MA ($Q_{top},Q_{bottom}$) 106948, 110284 218119, 210865 0.92, 0.865
EN ($Q_{top},Q_{bottom}$) 19609, 20672 21499, 21800 0.702, 0.71
CH ($Q_{top},Q_{bottom}$) 5350, 5336 5953, 6069 0.79, 0.73
MO ($Q_{top},Q_{bottom}$) 5093, 6041 5966, 6096 0.87, 0.69
Table 4: Comparison of network properties of $Q_{top}$ $Q_{bottom}$ AAGs.
Datasets Number of askers in $Q_{top}$ Number of answerer in $Q_{top}$ Number of asker in $Q_{bottom}$ Number of answerer in $Q_{bottom}$ overlap of askers overlap of answerers
SE 7697 2861 8933 4701 797 878
GD 5959 2249 6417 4941 843 1905
MA 89557 14072 88772 39909 6891 18397
EN 14533 4542 15564 6349 1148 1241
CH 4408 1180 4021 1860 427 545
MO 4497 762 4724 1789 256 472
Table 5: Asker and Answerer behaviour $Q_{top}$ $Q_{bottom}$
Cross platform users: We found around 2739 users who are present in all the six platforms. Fraction of common users posted at least one question is around 29%-30% for $MA$ and $EN$ and for $SE$ it is 18%. However for $CH$, $GD$ and $MO$ platforms, this fraction comparatively lower ($\leq 15\%$). The observations are similar for the number of common users posting an answer – the popular platforms have far larger fractions compared to less popular ones. Next we compute the number of common users who participated in asking/answering a question in $Q_{top}$ and $Q_{bottom}$. The percentage of common users asking one or more questions in case of $MA$, $EN$ and $SE$ are 19.82%, 17.30% and 9.71% respectively in $Q_{top}$ and 20.0%, 16.24% and 9.49% respectively in $Q_{bottom}$. The percentage of common users in these three platforms who answered one or more questions in $Q_{top}$ are 9.3%, 8.72% and 6.17% respectively. In $Q_{bottom}$ these percentages are 12.92%, 7.7% and 6.4% respectively. For $MO$, $GD$ and $CH$, the percentage of common users acting as asker in $Q_{t}op$ are 7.84%, 6.38% and 5.69% respectively and in $Q_{bottom}$ are 8.06%, 6.09% and 5.69% respectively. For these three platforms, the percentage of common users acting as answerer in $Q_{t}op$ are 2.7%, 4.3% and 1.8% respectively and in $Q_{bottom}$ are 4.2%, 4.9% and 2.1% respectively.
Engagement of common users: In order to investigate the involvement of common users (present in $Q_{top}$ and $Q_{bottom}$ as answerers) in the overall platform, we compute the average number of questions/answers posted by them at an aggregate level. For $MA$, $EN$ and $SE$, the average number of questions posted by common users in $Q_{top}$ are 22.68, 8.5 and 3.04 and that in $Q_{bottom}$ are 21.60, 8.74 and 3.57 respectively. For these three platforms, the average number of answers posted by common users in $Q_{top}$ are 228.83, 47.32 and 61.01 and that in $Q_{bottom}$ are 166.7, 52.20 and 56.74 respectively. For $MO$, $GD$ and $CH$, the average number of questions posted by common users in $Q_{top}$ are 11.56, 5.1 and 5.61 respectively. and in $Q_{bottom}$ are 9.02, 5.45 and 7.55 respectively. For these three platforms, the average number of answers posted by common users in $Q_{top}$ are 49.81, 24.74 and 15.55 respectively. The average number of answers posted by common users in $Q_{bottom}$ are 35.11, 21.43 and 14.22 respectively. Thus (i) irrespective of the popularity of the platforms, common users engage more in terms of asking and answering in $Q_{top}$ than in $Q_{bottom}$ and (ii) quite interestingly, the average number of answers posted by the common users in $Q_{top}$ and $Q_{bottom}$ are far more than the questions posted by them which indicates that these users are more engaged in posting answers than questions.
4 Question Category Prediction
4.1 Classification model
We consider previously discussed properties of the questions as features and predict whether a question will belong to $Q_{top}$ or $Q_{bottom}$. The features are – length of question body, Flesch reading ease, Coleman Liau index, automated readability index, number of tags, reputation of the asker and the clustering coefficient of asker in the asker-answerer graph. For every dataset, we have randomly divided the whole dataset into three parts – training (70%), validation (10%), and test (20%). Given a dataset, the number of instances present in training, validation, and test are given in Table 6. For this experiment, we have used four standard machine learning classifiers – (a) logistic regression (LR), (b) support vector classifier (SVC), (c) random forest (RF) (d) XGBoost. For evaluation, we have used overall precision, recall and macro F1 score.
Datasets Train Size Valid Size Test Size
${SE}$ 15848 2268 4525
${GD}$ 12660 1812 3615
${MA}$ 325153 46543 92809
${EN}$ 31119 4454 8884
${CH}$ 8844 1266 2525
${MO}$ 8875 1270 2534
Table 6: Number of instances present in training, validation and testing for all the datasets
Parameter settings: The parameter settings for each model are given below. For SVC, we consider the value of parameter $C$ in the range of 1-3. The values tried for kernel are poly, rbf and sigmoid. For random forest classifier, the range of the number of estimators is between 160 to 200. The criterion values are gini and entropy. For XGBoost, the range of values for the number of estimators is from 160 to 200. The range of learning rate is 0.1 to 0.3.
Logistic regression (LR): For all datasets, we found the default parameters are the best.
Support vector classifier (SVC): For $CH$, the best value of $C$ is 3, and kernel is set to poly. For other datasets, the best parameter for kernel is rbf. For $EN$ and $GD$, the value of the parameter C is set to 2 and 1 respectively. For the other datasets, the value of C is 3.
Random forest (RF): For $CH$, $EN$, $MO$, $MA$, the best value for criterion is entropy. For $GD$ and $SE$, the criterion is set to gini. For $MO$, the number of estimators is set to 180. For other datasets, the number of estimators is set to 200.
XGBoost: For all the datasets, the best value for the learning rate is 0.1. For $CH$, $GD$, $MO$, the value of the number of estimators is set to 160. For $EN$ and $SE$, the number of estimators is 180. For $MA$, the number of estimators is 200.
4.2 Results
Datasets LR SVC RF XGBoost
Precision Recall Macro F1 Precision Recall Macro F1 Precision Recall Macro F1 Precision Recall Macro F1
${SE}$ 0.67 0.67 0.67 0.69 0.69 0.69 0.71 0.71 0.71 0.72 0.72 0.72
${GD}$ 0.62 0.61 0.61 0.63 0.62 0.62 0.62 0.62 0.62 0.63 0.63 0.63
${MA}$ 0.69 0.69 0.69 0.70 0.70 0.70 0.69 0.69 0.69 0.71 0.71 0.71
${EN}$ 0.61 0.61 0.60 0.61 0.61 0.61 0.63 0.63 0.63 0.64 0.64 0.64
${CH}$ 0.59 0.58 0.58 0.59 0.58 0.57 0.59 0.58 0.59 0.58 0.58 0.58
${MO}$ 0.56 0.56 0.56 0.56 0.56 0.56 0.56 0.56 0.56 0.57 0.57 0.57
Table 7: Results of question classification models for all the datasets. The best results are highlighted in bold. The $2^{nd}$ best results are highlighted in underline.
In Table 7, we present the overall precision, recall and macro F1 score for all the datasets. For $SE$, RF and XGBoost are performing better than other models. For $GD$, almost all the models’ performances are similar (0.61-0.63 F1 score). For $MA$, XGBoost model is performing better (0.71 macro F1 score) than other models. For $EN$, RF and XGBoost attain better F1 score than other models. For $CH$, RF performs better (0.59 F1 score) than other models. For $MO$, all the models attain similar performance (0.56-0.57 F1 score).
5 Conclusion
In this paper, we study question-answering trends across six diverse CQA platforms. We find metadata, question structure, and user interaction patterns affect response time. Shorter, clearer questions with fewer tags are answered faster. High-reputation users engage more with quickly-answered questions. These question and asker features predict fast responses. We use machine learning to classify new questions based on metadata, planning to incorporate these factors into deep learning models to predict response times for new questions using text, metadata, and asker features.